Your company is looking at a new project in Mexico. The project will cost 900,000 pesos. The cash flows are expected to be 400,000 pesos per year for 5 years. The current spot exchange rate is 19.07 pesos per dollar. The risk-free rate in the US is 4%, and the risk-free rate in Mexico 8%. The dollar required return is 10%. What is the net present value of this investment in U.S. Dollars?

PLEASE SHOW IN EXCEL
Answer #1

Exchange Rate over 5 years:

1 USD = 19.07 Pesos

Exchange Rate after 1 Year = 19.07 * 1.08 / 1.04

= 19.80

Exchange Rate after 2 Year = 19.80 * 1.08 / 1.04

= 20.57

Exchange Rate after 3 Year = 20.57 * 1.08 / 1.04

= 21.36

Exchange Rate after 4 Year = 21.36 * 1.08 / 1.04

= 22.18

Exchange Rate after 5 Year = 22.18 * 1.08 / 1.04

= 23.03

FWd rate = Spot rate * [ 1 + Int Rate in Mexico ] / [ 1 + Int rate in US ]

NPV = PV of Cash Inflows - PV of Cash Outflows.

Year CF Exchange Rate CF in Dollar PVF @10% Disc CF
0 -900000 19.07 $ -47,194.55     1.0000 $ -47,194.55
1 400000 19.8 $   20,202.02     0.9091 $ 18,365.47
2 400000 20.57 $   19,445.79     0.8264 $ 16,070.90
3 400000 21.36 $   18,726.59     0.7513 $ 14,069.57
4 400000 22.18 $   18,034.27     0.6830 $ 12,317.65
5 400000 23.03 $   17,368.65     0.6209 $ 10,784.56
NPV $ 24,413.61